# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 1.2 Overview of Techniques
Prior Techniques and Obstacles for Learning Noisy Multi-Index Models.
Recall that our goal is, given access to noisy samples from well-behaved MIM with finite label set, to learn the underlying function $f$ to high accuracy. We will use $W$ to denote the hidden $K$-dimensional subspace to whose projection $f$ depends. The standard approach for solving such learning problems (see, e.g., [Vem10b]) is to first learn some suitable approximation $V$ to $W$, and then use exhaustive search to learn a function of the projection onto $V$. Computing such an approximation to $W$ is usually achieved by estimating the low-order moments of the level sets of $f$. In particular, for each output value $i$ and each low-degree polynomial $p$, we want to approximate the expectation of $p(\mathbf{x})\mathds{1}(f(\mathbf{x})=i)$. In the absence of label noise, these moments should depend solely on how $p$ interacts with the directions aligned with $W$. Using this property, one can extract an approximation of a subspace of $W$, by choosing $V$ to be the space spanned by all directions with reasonably large moments. This approach yields an algorithm with runtime approximately $d^{O(m)}$, where $m$ is the number of moments being computed, plus some (often exponential) function of $\dim(V)$—which is typically regarded as “constant”, since $d$ is much larger than $\dim(V)$.
@@
New Dimensionality Reduction Technique.
@@
One disadvantage of our method as described above is that, while the runtime is typically polynomial in the dimension $d$, the dependence on other parameters can be problematic. Specifically, the number of regions $S$ is usually exponential in $\dim(V)$; hence, at each iteration, $\dim(V)$ can increase in the worst-case by an exponential amount. This can lead to power-tower dependencies on some parameters. We circumvent this issue by assuming that our function has a small fraction of regions $S$ with non-vanishing moments. With this assumption, we can only add directions with large moments in a reasonable fraction of regions using a filtering step (see Line 6), keeping the increase in $\dim(V)$ polynomial. This bound can be further improved if the noisy version of $f(\mathbf{x})$ that we observe is actually a function of the projection onto $W$, which holds in both the realizable and the RCN settings (see Section 4.2). In this case, non-vanishing moments will exactly align with directions along $W$, up to a measurement error, ensuring that $\dim(V)$ never exceeds $\dim(W)$ (see Theorem 4.9).
SQ Lower Bounds.
@@
Applications to Well-studied Concept Classes.
@@ 1.3 Prior Work
The recent work [DKK+24] developed agnostic learners for MIMs under the Gaussian distribution with fixed-parameter tractability guarantees similar to ours. The major difference is that, while our framework relies on learning from random examples, the latter work [DKK+24] assumes query access to the target function—a much stronger access model.
@@ 1.4 Notation and Preliminaries
Basic Notation For $n\in\mathbb{Z}_{+}$, let $[n]:=\{1,\ldots,n\}$. We use small boldface characters for vectors and capital bold characters for matrices. For $\mathbf{x}\in\mathbb{R}^{d}$ and $i\in[d]$, $\mathbf{x}_{i}$ denotes the $i$-th coordinate of $\mathbf{x}$, and $\|\mathbf{x}\|_{2}:=(\mathop{\textstyle\sum}_{i=1}^{d}\mathbf{x}_{i}^{2})^{1/2}$ denotes the $\ell_{2}$-norm of $\mathbf{x}$. Throughout this text, we will often omit the subscript and simply write $\|\mathbf{x}\|$ for the $\ell_{2}$-norm of $\mathbf{x}$. We will use $\mathbf{x}\cdot\mathbf{y}$ for the inner product of $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ and $\theta(\mathbf{x},\mathbf{y})$ for the angle between $\mathbf{x},\mathbf{y}$. For vectors $\mathbf{x},\mathbf{v}\in\mathbb{R}^{d}$ and a subspace $V\subseteq\mathbb{R}^{d}$ denote by $\mathbf{x}^{V}$ the projection of $\mathbf{x}$ onto $V$, and by $\mathbf{x}^{\mathbf{v}}$ the projection of $\mathbf{x}$ onto the line spanned by $\mathbf{v}$. For two subspaces $V,W\subseteq\mathbb{R}^{d}$, we denote by $W^{V}=\{\mathbf{w}^{V}:\mathbf{w}\in W\}$, which it itself a subspace. Furthermore, for a set of vectors $L\subseteq\mathbb{R}^{d}$, we denote by $\mathrm{span}(L)$ the subspace of $\mathbb{R}^{d}$ defined by their span. We slightly abuse notation and denote by $\mathbf{e}_{i}$ the $i$-th standard basis vector in $\mathbb{R}^{d}$. For a matrix $A\in\mathbb{R}^{m\times n}$, the Frobenius norm of $A$ is denoted by $\left\|A\right\|_{F}=(\sum_{i=1}^{m}\sum_{j=1}^{n}A_{i,j}^{2})^{1/2}$. For a set $A\subseteq\mathbb{R}^{d}$ and a point $\mathbf{x}\in\mathbb{R}^{d}$, we define the distance from $\mathbf{x}$ to $A$ as $\mathrm{dist}(\mathbf{x},A)=\min_{\mathbf{y}\in A}\left\|\mathbf{x}-\mathbf{y}\right\|$.
@@ 1.5 Organization
The paper is structured as follows: In Section 2, we present a special case of our subspace approximation procedure for learning multiclass linear classifiers under agnostic noise, establishing Theorem 1.6. In Section 3, we generalize this approach, stating the necessary conditions (see Definition 1.2) for our algorithm to work for general classes of MIMs, establishing Theorem 1.4. In Section 4, we consider applications of our algorithmic approach. First in Section 4.1, we apply our general theorem to the class of intersections of halfspaces, establishing Theorem 1.9. Moreover, in section Section 4.2, we consider how our algorithm performs under random noise, particularly in the Random Classification Noise (RCN) model, and demonstrate that we can get even better runtime guarantees (Theorem 1.8). Finally, Section 5 provides SQ lower bounds. First, we demonstrate that the complexity of our algorithm is qualitatively optimal with respect to the dimension, establishing Theorem 1.5. Second, we establish an SQ lower bound for learning multiclass linear classifiers under random classification noise, showing that achieving optimal error with a polynomial number of samples is impossible in the SQ model, Theorem 1.7.
@@ 2 Learning Multiclass Linear Classifiers with Adversarial Label Noise
Let $\mathcal{L}_{d,K}$ be the class of $K$-class linear classifiers in $\mathbb{R}^{d}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $\mathrm{OPT}=\inf_{f\in\mathcal{L}_{d,K}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$ and let $\epsilon,\delta\in(0,1)$. There exists a sufficiently large universal constant $C>0$ such that Algorithm 1 draws $N=d2^{\mathrm{poly}(1/\epsilon,K)}\log(1/\delta)$ i.i.d. samples from $D$, runs in $\mathrm{poly}(d,N)$ time, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq{C}\cdot\mathrm{OPT}+\epsilon$.
@@
Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ with an orthonormal basis $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$, and let $\epsilon\in(0,1)$. We call an $\epsilon$-approximating partition with respect to $V$ a family of sets $\mathcal{S}$ of the form $S_{\mathbf{j}}=\{\mathbf{x}\in\mathbb{R}^{d}:z_{\mathbf{j}_{i}-1}\leq\mathbf{v}^{(i)}\cdot\mathbf{x}\leq z_{\mathbf{j}_{i}},i\in[k]\}$, where $\mathbf{j}\in[M_{\epsilon}]^{k}$, $z_{i}=-\sqrt{2\log(k/\epsilon)}+i\epsilon+t$, for $i\in[M_{\epsilon}]$, $t\in(0,\epsilon/2)$ and $M_{\epsilon}=\left\lceil(2\sqrt{2\log(k/\epsilon)})/\epsilon\right\rceil$. Furthermore, for each set $S_{\mathbf{j}}\in\mathcal{S}$, we associate a set of linear equation $\mathcal{G}_{S_{\mathbf{j}}}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=t_{1},\dots,\mathbf{v}^{(k)}\cdot\mathbf{x}=t_{k}\}$, where $z_{\mathbf{j}_{i}-1}\leq t_{i}\leq z_{\mathbf{j}_{i}}$ for $i\in[k]$.
Furthermore, we define the piecewise constant approximation of a distribution as the function that minimizes the 0-1 loss with respect to the labels in each subset of a partition. Our algorithm will first estimate the hidden subspace of the multiclass linear classifier and then return a piecewise constant approximation over it.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$, let $V$ be a subspace of $\mathbb{R}^{d}$, let $\epsilon\in(0,1)$, and let $\mathcal{S}$ be an $\epsilon$-approximating partition with respect to $V$. We call a piecewise constant approximation of the distribution $D$, with respect to a partition $\mathcal{S}$, any classifier $h_{{\mathcal{S}}}:\mathbb{R^{d}}\to[K]$ of the form $h_{{\mathcal{S}}}(\mathbf{x})=\operatorname*{argmin}_{i\in[K]}(\operatorname*{\mathbf{Pr}}_{{(\mathbf{x},y)\sim D}}[i\neq y\mid\mathbf{x}\in S])$, if $\mathbf{x}\in S$ for all $S\in\mathcal{S}$, where $\mathcal{S}$ is an $\epsilon$-approximating partition with respect to $V$ according to Definition 2.2.
We now present our subspace approximation method for learning multiclass linear classifiers under agnostic noise, Algorithm 1.
Input: Accuracy $\epsilon>0$, failure probability $\delta>0$ sample access to a distribution $D$ over $\mathbb{R}^{d}\times[K]$.
Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}{+}\epsilon$, for a universal constant $C$.
@@
    Let $c$ be a sufficiently large constant, and let $T{,p}$ be sufficiently large, constant-degree polynomials of $1/\epsilon$ and $K$.
@@
      Draw a set $S$ of $N$ i.i.d. samples from $D$.
@@
      $\mathcal{E}_{t}\leftarrow$ Algorithm 2($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
@@
    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as described in Definition 2.3, using $N$ i.i.d. samples from $D$, by finding the most frequent label at any set in $\mathcal{S}$.
@@
Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$, and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
    Construct an $\epsilon^{\prime}$-approximating partition, $\mathcal{S}$, with respect to $V$ (see Definition 2.2).
@@
    Compute: $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$ for $S\in\mathcal{S},i\in[K]$.
@@
Let $f^{*}$ be a multiclass linear classifier with error $\mathrm{OPT}$, i.e., $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$. Algorithm 1 operates as follows: In each iteration $t$, the algorithm updates the current list of unit vectors $L_{t}$ so that its span, denoted by $V_{t}$, provides a better approximation to the span of the weight vectors of $f^{*}$, denoted by $W^{*}$. To this end, Algorithm 1 uses Algorithm 2 to construct a new list $\mathcal{E}_{t}$ that contains vectors with non-trivial correlation to $(W^{*})^{\perp V_{t}}$. Our main contribution is the design and analysis of Algorithm 2, which efficiently returns a non-trivially correlating vector.
@@
First in Section 2.1, we show that if an optimal multiclass classifier is not $C\cdot\mathrm{OPT}$-close to a constant classifier, then at least one of the first moments of the labels regions, i.e., $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$ for some $i\in[K]$, will correlate non-trivially (by at least $\mathrm{poly}(\epsilon)$) with a vector in $W^{*}$. This, in fact, suffices to initialize a list of vectors that correlate with $W^{*}$. In order to expand this list and at the same time improve the approximation of $W^{*}$, we need to show that, we can find vectors that not only correlate with $W^{*}$ but are also orthogonal to the current list of vectors $L_{t}$.
@@
To this end, in Section 2.2, we use the current list $L_{t}$ of weight vectors to partition $\mathbb{R}^{d}$ into sufficiently small regions $S$. In each such region $S$, the classifier $f^{*}$ has a limited dependence on the vectors in $L_{t}$. Specifically, we show that $f^{*}$ is $\epsilon$-close to a function $f_{S}^{*}$ that does not depend on weight vectors contained in $L_{t}$. Thus, within each region $S$, we obtain a multiclass classifier that depends only on vectors in $(W^{*})^{\perp V_{t}}$. Therefore, from the above discussion, in each region $S$, either the classifier can be accurately approximated by a constant classifier, or the first moments will correlate with a vector in $(W^{*})^{\perp V_{t}}$, enabling further improvement.
@@
Furthermore, for $f^{*}$ to depend mostly on directions in $(W^{*})^{\perp V_{t}}$, the cubes $S$ in the partition must have width on the order of $\mathrm{poly}(\epsilon/K)$. Such a fine partitioning would typically yield a number of cubes that is exponential in $\dim(V_{t})$ and, hence could lead to a power-tower update in the dimension of $V_{t}$ when updating $L_{t}$. In order to avoid this power-tower dependence, we leverage the fact that the total probability mass of cubes with non-trivial correlation is at least $\mathrm{poly}(\epsilon/K)$ (this will be proved in the analysis of Algorithm 2 in Proposition 2.8). By weighting each moment vector by the probability mass of its corresponding cube (see Line 5), we are guaranteed to find a vector with non-trivial correlation by examining only a $\mathrm{poly}(K/\epsilon)$-sized list of candidate vectors.
@@
We note that our algorithm is not proper; that is, it does not return a multiclass linear classifier. Instead, it produces a piecewise constant function over a sufficiently fine partition of $\mathbb{R}^{d}$.

### v2

@@ 1.2 Overview of Techniques
Prior Techniques and Obstacles for Learning Multi-Index Models
Recall that our goal is, given access to noisy samples from a well-behaved MIM with finite label set, to learn the underlying function $f$ to high accuracy. We will use $W$ to denote the hidden $K$-dimensional subspace to whose projection $f$ depends. The standard approach for solving such learning problems (see, e.g., [Vem10b]) is to first learn some suitable approximation $V$ to $W$, and then use exhaustive search to learn a function of the projection onto $V$. Computing such an approximation to $W$ is usually achieved by estimating the low-order moments of the level sets of $f$. In particular, for each output value $i$ and each low-degree polynomial $p$, we want to approximate the expectation of $p(\mathbf{x})\mathds{1}(f(\mathbf{x})=i)$. In the absence of label noise, these moments should depend solely on how $p$ interacts with the directions aligned with $W$. Using this property, one can extract an approximation of a subspace of $W$, by choosing $V$ to be the space spanned by all directions with reasonably large moments. This approach yields an algorithm with runtime approximately $d^{O(m)}$, where $m$ is the number of moments being computed, plus some (often exponential) function of $\dim(V)$—which is typically regarded as “constant”, since $d$ is much larger than $\dim(V)$.
@@
New Dimensionality Reduction Technique
@@
This generalization of the previously used one-shot dimension reduction technique has two main advantages. Practically speaking, the more relaxed condition is often easier to prove. This allows, for example, our qualitative improvements in the error for robustly learning intersections of halfspaces. As far as we know, the one-shot dimension reduction technique may suffice to obtain error $\tilde{O}(\mathrm{OPT})$ in this case. However, the difficulty of proving such results means that the best analysis we know of gives substantially worse bounds (and only after a quite involved analysis). On the other hand, the condition needed for our iterative dimension reduction is somewhat easier to obtain, giving us a substantially better $\tilde{O}(K\mathrm{OPT})$ error with a simpler analysis.
Furthermore, our new technique allows us to efficiently learn a broader family of MIMs compared to what was previously possible with one-shot dimension reduction techniques, even in principle. In particular, as we show in Appendix F, there exist Multi-Index Models that cannot be learned to better than constant error with any constant number of moments using the one-shot technique; and on the other hand, can be efficiently learned to arbitrary accuracy using only second moments via our iterative technique.
One disadvantage of our method as described above is that, while the runtime is typically polynomial in the dimension $d$, the dependence on other parameters can be problematic. Specifically, the number of regions $S$ is usually exponential in $\dim(V)$; hence, at each iteration, $\dim(V)$ can increase in the worst-case by an exponential amount. This can lead to power-tower dependencies on some parameters. We circumvent this issue by assuming that our function has a small fraction of regions $S$ with non-vanishing moments. With this assumption, we can only add directions with large moments in a reasonable fraction of regions using a filtering step (see Line 6), keeping the increase in $\dim(V)$ polynomial. This bound can be further improved if the noisy version of $f(\mathbf{x})$ that we observe is actually a function of the projection onto $W$, which holds in both the realizable and the RCN settings (see Section 4.2). In this case, non-vanishing moments will exactly align with directions along $W$, up to a measurement error, ensuring that $\dim(V)$ never exceeds $\dim(W)$ (see Theorem 4.9). For an illustration of our iterative procedure for the special case of multiclass linear classification, see Figure 1 in Section 2.
SQ Lower Bounds
@@
Applications to Well-studied Concept Classes
@@ 1.3 Prior Work
The recent work [DKK+23] developed agnostic learners for MIMs under the Gaussian distribution with fixed-parameter tractability guarantees similar to ours. The major difference is that, while our framework relies on learning from random examples, the latter work [DKK+23] assumes query access to the target function—a much stronger access model.
@@ 1.4 Notation and Preliminaries
Basic Notation For $n\in\mathbb{Z}_{+}$, let $[n]:=\{1,\ldots,n\}$ and $\overline{\mathbb{R}}\coloneqq\mathbb{R}\cup\{\pm\infty\}$. We denote by $2^{S}=\{T\mid T\subseteq S\}$ the power set of the set $S$. We use small boldface characters for vectors and capital bold characters for matrices. For $\mathbf{x}\in\mathbb{R}^{d}$ and $i\in[d]$, $\mathbf{x}_{i}$ denotes the $i$-th coordinate of $\mathbf{x}$, and $\|\mathbf{x}\|_{2}:=(\mathop{\textstyle\sum}_{i=1}^{d}\mathbf{x}_{i}^{2})^{1/2}$ denotes the $\ell_{2}$-norm of $\mathbf{x}$. Throughout this text, we will often omit the subscript and simply write $\|\mathbf{x}\|$ for the $\ell_{2}$-norm of $\mathbf{x}$. We will use $\mathbf{x}\cdot\mathbf{y}$ for the inner product of $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ and $\theta(\mathbf{x},\mathbf{y})$ for the angle between $\mathbf{x},\mathbf{y}$. For vectors $\mathbf{x},\mathbf{v}\in\mathbb{R}^{d}$ and a subspace $V\subseteq\mathbb{R}^{d}$ denote by $\mathbf{x}^{V}$ the projection of $\mathbf{x}$ onto $V$, $\mathbf{x}^{\perp V}$ the projection of $\mathbf{x}$ onto the orthogonal complement of $V$ and by $\mathbf{x}^{\mathbf{v}}$ the projection of $\mathbf{x}$ onto the line spanned by $\mathbf{v}$. For two subspaces $V,W\subseteq\mathbb{R}^{d}$, we denote by $W^{V}=\{\mathbf{w}^{V}:\mathbf{w}\in W\}$ and by $V+W=\{\mathbf{w}+\mathbf{v}:\mathbf{w}\in W,\mathbf{v}\in V\}$, note that both $W^{V}$ and $V+W$ are subspaces. Furthermore, for a set of vectors $L\subseteq\mathbb{R}^{d}$, we denote by $\mathrm{span}(L)$ the subspace of $\mathbb{R}^{d}$ defined by their span. We slightly abuse notation and denote by $\mathbf{e}_{i}$ the $i$-th standard basis vector in $\mathbb{R}^{d}$. For a matrix $A\in\mathbb{R}^{m\times n}$, the Frobenius norm of $A$ is denoted by $\left\|A\right\|_{F}=(\sum_{i=1}^{m}\sum_{j=1}^{n}A_{i,j}^{2})^{1/2}$. For a set $A\subseteq\mathbb{R}^{d}$ and a point $\mathbf{x}\in\mathbb{R}^{d}$, we define the distance from $\mathbf{x}$ to $A$ as $\mathrm{dist}(\mathbf{x},A)=\min_{\mathbf{y}\in A}\left\|\mathbf{x}-\mathbf{y}\right\|$.
@@ 1.5 Organization
The paper is structured as follows: In Section 2, we present a special case of our subspace approximation procedure for learning multiclass linear classifiers under agnostic noise, establishing Theorem 1.8. In Section 3, we generalize this approach, stating the necessary conditions (see Definition 1.2) for our algorithm to work for general classes of MIMs, establishing Theorem 1.5. In Section 4, we present additional applications of our algorithmic approach. In Section 4.1, we apply our general theorem to the class of intersections of halfspaces, establishing Theorem 1.13. In Section 4.2, we adapt our general algorithm for the case of random noise, and demonstrate that it yields qualitative stronger complexity guarantees (Theorem 1.11). Finally, Section 5 gives our SQ lower bounds. First, we demonstrate that the complexity of our algorithm is qualitatively optimal as a function of the dimension, establishing Theorem 1.6. Second, we give an SQ lower bound (Theorem 1.9) for learning multiclass linear classifiers under random classification noise (and in the presence of coarse labels).
@@ 2 Learning Multiclass Linear Classifiers with Adversarial Label Noise
Let $\mathcal{L}_{d,K}$ be the class of $K$-class linear classifiers in $\mathbb{R}^{d}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $\mathrm{OPT}=\inf_{f\in\mathcal{L}_{d,K}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$ and let $\epsilon,\delta\in(0,1)$. There exists a sufficiently large universal constant $C>0$ such that Algorithm 1 draws $N=d2^{\mathrm{poly}(1/\epsilon,K)}\log(1/\delta)$ i.i.d. samples from $D$, runs in $\mathrm{poly}(N)$ time, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$.
@@
Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ with an orthonormal basis $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$, and let $\epsilon\in(0,1)$. An $\epsilon$-approximating partition with respect to $V$ is a collection of sets $\mathcal{S}$ defined as follows: For each multi-index $\mathbf{j}=(\mathbf{j}_{1},\dots,\mathbf{j}_{k})\in[M_{\epsilon}]^{k}$, define $S_{\mathbf{j}}=\{\mathbf{x}\in\mathbb{R}^{d}:z_{\mathbf{j}_{i}-1}\leq\mathbf{v}^{(i)}\cdot\mathbf{x}\leq z_{\mathbf{j}_{i}},i\in[k]\}$ where $z_{i}$’s are defined as $z_{i}=-\sqrt{2\log(k/\epsilon)}+i\epsilon+t$, for $i\in\{0,\dots,M_{\epsilon}\},t\in(0,\epsilon/2)$ and $M_{\epsilon}=\left\lceil(2\sqrt{2\log(k/\epsilon)})/\epsilon\right\rceil$. Furthermore, to each set $S_{\mathbf{j}}\in\mathcal{S}$, we associate a set of linear equations $\mathcal{G}_{S_{\mathbf{j}}}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=t_{1},\dots,\mathbf{v}^{(k)}\cdot\mathbf{x}=t_{k}\}$, where for each $i\in[k]$, $t_{i}$ is chosen so that $z_{\mathbf{j}_{i}-1}\leq t_{i}\leq z_{\mathbf{j}_{i}}$.
Furthermore, we define the piecewise constant approximation of a distribution over labeled examples as the function that minimizes the 0-1 loss with respect to the labels in each subset of a partition. Our algorithm will first estimate the hidden subspace of the multiclass linear classifier and then return a piecewise constant approximation over it.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$, let $V$ be a subspace of $\mathbb{R}^{d}$ and let $\epsilon\in(0,1)$. Let $\mathcal{S}$ be an $\epsilon$-approximating partition with respect to $V$ (see Definition 2.2). A piecewise constant approximation of the distribution $D$, with respect to $\mathcal{S}$, is any classifier $h_{\mathcal{S}}:\mathbb{R}^{d}\to[K]$ such that for each $S\in\mathcal{S}$ and $\mathbf{x}\in S$, $h_{\mathcal{S}}$ is defined as follows $h_{\mathcal{S}}(\mathbf{x})\in\operatorname*{argmin}_{i\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[i\neq y\mid\mathbf{x}\in S]$.
We now present our subspace approximation method for learning multiclass linear classifiers under agnostic noise (see Algorithm 1).
Input: Accuracy $\epsilon>0$, failure probability $\delta>0$ sample access to a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}{+}\epsilon$, where $C$ is a universal constant and $\mathrm{OPT}\coloneqq\inf_{f\in\mathcal{L}_{d,K}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
@@
    Let $c$ be a sufficiently large constant, and let $T$ and $p$ be sufficiently large, constant-degree polynomials in $1/\epsilon$ and $K$.
@@
      Draw a set $S_{t}$ of $N$ i.i.d. samples from $D$.
@@
      $\mathcal{E}_{t}\leftarrow$ Algorithm 2($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S_{t}$).
@@
    Construct $\mathcal{S}$, an $\epsilon^{\prime}$-approximating partition with respect to $\mathrm{span}(L_{t})$ (see Definition 2.2).
@@
    Draw $N$ i.i.d. samples from $D$ and construct the piecewise constant classifier $h_{\mathcal{S}}$ as follows: For each $S\in\mathcal{S}$, assign a label that appears most frequently among the samples falling in $S$. Return $h_{\mathcal{S}}$.
@@
Input: $\epsilon,\epsilon^{\prime},\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$, and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
    Construct an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ with respect to $V$ (see Definition 2.2).
@@
    For each $S\in\mathcal{S}$ and for each label $i\in[K]$ compute $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$.
@@
Let $f^{*}$ be a $K$-multiclass linear classifier with error $\mathrm{OPT}$, i.e., $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$. Algorithm 1 proceeds in a sequence of iterations: In each iteration $t$, it maintains and updates a set of unit vectors $L_{t}$ so that its span $V_{t}$ provides a better approximation to the span of the weight vectors of $f^{*}$, denoted by $W^{*}$. To this end, Algorithm 1 uses Algorithm 2 to construct a new list $\mathcal{E}_{t}$ that contains vectors with non-trivial correlation to $(W^{*})^{\perp V_{t}}$.
@@
First, in Section 2.1, we show that if $f^{*}$ is not $C\cdot\mathrm{OPT}$-close to any constant classifier, then there exists at least one label $i\in[K]$ for which the first moment $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$ correlates non-trivially (by at least $\mathrm{poly}(\epsilon/K)$) with a vector in $W^{*}$. This, in fact, suffices to initialize a list of vectors that correlate with a vector in $W^{*}$. In order to expand this list and at the same time improve the approximation of $W^{*}$, we need to show that, we can find vectors that not only correlate with $W^{*}$ but are also orthogonal to the current list of vectors $L_{t}$.
@@
To this end, Algorithm 2 uses the current set $L_{t}$ of weight vectors to partition $\mathbb{R}^{d}$ into sufficiently small regions $S$. In each such region $S$, the classifier $f^{*}$ has a limited dependence on the vectors in $L_{t}$. Specifically, we show that $f^{*}$ is $\epsilon$-close to a function $f_{S}^{*}$ that does not depend on weight vectors contained in $L_{t}$. Thus, within each region $S$, we obtain a multiclass classifier that depends only on vectors in $(W^{*})^{\perp V_{t}}$. Therefore, in every region $S$, either the classifier can be well-approximated by a constant function or the first moments will correlate non-trivially with a vector in $(W^{*})^{\perp V_{t}}$, thus enabling further improvement of the approximation.
@@
Furthermore, for $f^{*}$ to depend mostly on directions in $(W^{*})^{\perp V_{t}}$, the cubes $S$ in the partition must have width on the order of $\mathrm{poly}(\epsilon/K)$. Such a fine partitioning would typically yield a number of cubes that is exponential in $\dim(V_{t})$ and, hence could lead to a power-tower update in the dimension of $V_{t}$ when updating $L_{t}$. In order to avoid this power-tower dependence, we leverage the fact that the total probability mass of cubes with non-trivial correlation is at least $\mathrm{poly}(\epsilon/K)$ (this will be proved in the analysis of Algorithm 2 in Proposition 2.8). By weighting each moment vector by the probability mass of its corresponding cube (see Line 5 of Algorithm 2), we are guaranteed to find a vector with non-trivial correlation by examining only a $\mathrm{poly}(K/\epsilon)$-sized list of candidate vectors.
@@
As follows from the above, our algorithm is not proper; that is, it does not return a multiclass linear classifier. Instead, it produces a piecewise constant function over a sufficiently fine partition of $\mathbb{R}^{d}$.
See Figure 1 for an illustration.
Organization of Section 2